40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed May 15 00:40:19 2019, gautam, Update, SUS, ETMY diagnosis plan ETMY_sensorSpectra.pdf
    Reply  Wed May 15 10:57:57 2019, gautam, Update, SUS, EY chamber opened 
       Reply  Wed May 15 17:46:24 2019, gautam, Update, SUS, ETMY inspection firstContactFiber.JPGETMY_sensorSpectra.pdf
          Reply  Wed May 15 19:36:29 2019, Koji, Update, SUS, ETMY instepction 
             Reply  Thu May 16 13:07:14 2019, gautam, Update, SUS, First contact residue removal etmy_sensors.pdfetmy_BRmode.pdf
                Reply  Thu May 16 23:31:55 2019, gautam, Update, SUS, ETMY suspension characterization 6x
                   Reply  Fri May 17 17:01:08 2019, gautam, Update, SUS, ETMY suspension characterization etmy_oplevs.pdf
          Reply  Mon May 20 11:33:46 2019, gautam, Update, SUS, ITMY inspection 
             Reply  Mon May 20 17:12:57 2019, gautam, Update, SUS, ETMY LL adjustment 
                Reply  Mon May 20 22:06:07 2019, gautam, Update, SUS, ITMY also kicked 
                   Reply  Tue May 21 00:15:21 2019, gautam, Update, SUS, Main objectives of vent achieved (?) etmy_oplevs_20190520.pdfETMY_cplxTF.pdfETMY_diagComp.pdfScreen_Shot_2019-05-21_at_12.37.08_AM.png
                      Reply  Tue May 21 21:33:27 2019, gautam, Update, SUS, ETMY HR face cleaned 
                         Reply  Wed May 22 11:53:50 2019, gautam, Update, SUS, ETMY EQ stops backed out ETMY_eigenmodes.pdf
Message ID: 14630     Entry time: Wed May 22 11:53:50 2019     In reply to: 14629
Author: gautam 
Type: Update 
Category: SUS 
Subject: ETMY EQ stops backed out 

Yesterday we noticed that the POS and SIDE eigenmodes were degenerate (with 1mHz spectral resolution). Moreover, the YAW peak had shifted down by ~500 mHz compared to earlier this week, although there was still good separation between PIT and YAW in the Oplev error signals. Ideas were (i) check if EQ stops were not backed out sufficiently, and (ii) look for any fibers/other constraints in the system. Today morning, I inspected the optic again. I felt the EQ stop viton tips were a bit close to the optic, so I backed them out further. Apart from this, I adjusted the LR and SIDE OSEM position in their respective holders to make the sensor voltages closer to half-light. Kicked the optic again just now, let's see if there is any change.

Remaining tasks:

  1. Check EY table leveling.
  2. Check EY actuation matrix diagonality using this technique.
  3. Check that IR resonances are seen (and all the usual pre-pumpdown alignment checks).
  4. Take close out pictures.
  5. Heavy doors on, pump down.

If everything goes smoothly, I think we should plan for the heavy doors going back on and commencing the pumpdown tomorrow. After discussion with Koji, we came to the conclusion that it isn't necessary to investigate IPANG (high likelihood of it falling off the steering optics during the pumpdown) / AS beam clipping (no strong evidence that this is a problem) for this vent.

Update 1235: Indeed, the eigenmodes are back to their positions from earlier this week. Indeed, the POS and SIDE modes are actually better separated! So, the OSEM/magnet and EQstop/optic interactions are non-negligible in the analysis of the dynamics of the pendulum.

Attachment 1: ETMY_eigenmodes.pdf  10 kB  Uploaded Wed May 22 13:41:58 2019  | Hide | Hide all
ETMY_eigenmodes.pdf
ELOG V3.1.3-